# Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads

### v1

@@ C.1 Combination with Quantization
C.2 Combination with LoCoCo
@@
Appendix D Training Robustness
@@
D.1 Intermediate Size of the retaining head
@@
D.2 Training Data Insensitivity
We also consider the sensitivity of the training data, which leads us to ablate the training dataset by training on LongAlign (Bai et al., 2024) and Anti-Haystack (Pan, 2024), comparing these results with those from LongAlpaca (Chen et al., 2024) in the original training setting. We also align other settings to the original setting and choose the backbone model to be Phi-3-mini-128K. We report the average L-Eval score for each training dataset. The results in Table 2 shows that Locret has high insensitivity towards different training data. The performance impact of different data recipes is minimal, indicating that our method can be trained on any long-context tuning dataset.
@@ D.2 Training Data Insensitivity
Appendix E Discontinuous Context and Stablizers
@@
Appendix F Causal Importance Score Simulates a Cache Problem
@@
Appendix G Retained Patterns of Locret
@@
Figure 6 shows that the pattern is mostly decided by the tasks, where both heads shows similar pattern in the same task. In R.Number, we are able to observe a strong signal between token 10000 and 15000, which is the position of the inserted number string, indicating that Locret can identify the potentially answer-related parts by giving high predicted values of CIS. In M.Find, we can observe the StreamingLLM (Xiao et al., 2024b) pattern, where the tokens at the beginning of the sequence are always important. This is also mentioned as the $\Lambda$-pattern in MInference. We can also discover the vertical lines in the middle of the sequence. This pattern is also approached by MInference (Jiang et al., 2024a) by the pattern “vertical-and-slash”. In E.MC, H2(Zhang et al., 2024e) and ScissorHands (Liu et al., 2024a) pattern can be observed, following the assumption that if a token is activated at some point, it will continue to be activated in the consequencing process. Noticing that the vertical lines always come in groups, which is the fundament of InfLLM (Xiao et al., 2024a) retrieving blocks to calculate. The comparison between two heads also shows that different heads exhibits different features. Head 22 of layer 11 shows stronger vertical lines at some point, where retained pattern of head 14 layer 11 is more even. Head 14 of layer 11 also gives stronger signal to the initial tokens, where this effect is less strong in head 22 layer 11. We also conduct experiments to investigate the patterns across layers. In Figure 6, we show that the pattern variance of the same head in different layers can be large. In shallow layers, e.g. layer 1 and 5, the retained cache units appears to be periodical and semantic independent. However, in middle layers, e.g. layer 13 and 17, the position of the inserted number string is strongly highlighted, indicating that semantic takes over to be the dominant factor. In the deepest layers, e.g. 21, 25 and 29, the highlighted vertical line at the position of the inserted string becomes more accurate.
@@
Appendix H The Training Process of Locret

### v2

@@ H.1 Combination with Quantization
H.2 Combination with LoCoCo
@@
Appendix I Training Robustness
Table 15: L-Eval scores of Locret trained on various datasets.
Dataset LongAlpaca LongAlign Anti-Haystack
L-Eval 21.33 22.00 20.72
@@ Appendix I Training Robustness
I.1 Intermediate Size of the retaining head
@@
I.2 Training Data Insensitivity
We also consider the sensitivity of the training data, which leads us to ablate the training dataset by training on LongAlign (Bai et al., 2024a) and Anti-Haystack (Pan, 2024), comparing these results with those from LongAlpaca (Chen et al., 2024) in the original training setting. We also align other settings to the original setting and choose the backbone model to be Phi-3-mini-128K. We report the average L-Eval score for each training dataset. The results in Table 6 shows that Locret has high insensitivity towards different training data. The performance impact of different data recipes is minimal, indicating that our method can be trained on any long-context tuning dataset.
@@ I.2 Training Data Insensitivity
Appendix J Extremely Long Context Evaluation
[…]
Table 18: Inference speed with Retaining Heads.
Phi-3-mini-128K on 10M context
Dataset R.PassKey_10M
Locret 100.00
Appendix K Compressing Multi-turn Conversations
Compared to query-aware eviction methods, such as SnapKV (Li et al., 2024a), Locret is a more suitable solution for multi-turn conversation scenarios. This is because the evaluation of cache importance in Locret is based on the cache itself, rather than being dependent on the subsequent query. To demonstrate this, we evaluate Locret on the Rock-Paper-Scissors benchmark introduced in SirLLM (Yao et al., 2024). Since SirLLM is specifically designed for such scenarios, we use it as our baseline in this benchmark. Results in Table 19 show that Locret is also effective in multi-turn conversation contexts.
The hyperparameters are aligned with those used in SirLLM, with the cache budget set to 1024, and no stabilizers are retained, as SirLLM does not retain local tokens in this benchmark. We perform 2000 turns as same as the original SirLLM settings. The results are presented below.
Table 19: Rock-Paper-Scissors scores of Locret and SirLLM.
Phi-3-mini-128K on Rock-Paper-Scissors
Preference Rock Paper Scissors Avg.
win tie lose win tie lose win tie lose $\uparrow$ $\downarrow$
SirLLM 40.00 31.75 28.25 27.5 36.55 35.96 29.35 25.15 45.50 32.28 36.57
Locret 18.95 50.00 31.05 30.35 19.45 50.20 52.05 27.25 20.70 33.78 33.98
Appendix L Discontinuous Context and Stablizers
@@
Appendix M Retaining Heads Do not Slow Down Inference
We evaluate the model’s forward throughput under varying context lengths, both with and without retaining heads. The results are summarized below in Table 20. “$\mathbf{R}$” represents the retaining heads, and the throughput is reported in tokens per second (tok/s) in the format “Avg. / Std.”
Table 20: Inference speed with Retaining Heads.
Context Length 1024 2048 3072 4096
w/o $\mathbf{R}$ Speed 18674 / 443 19743 / 464 19982 / 402 20304 / 187
w/ $\mathbf{R}$ Speed 17118 / 1117 18503 / 546 19054 / 283 19153 / 174
From the results, no significant latency increase is observed when using retaining heads. The numerical differences are attributed to systematic variations rather than additional overhead introduced by retaining heads during inference.
Appendix N Causal Importance Score Simulates a Cache Problem
@@
Appendix O Retained Patterns of Locret
@@
Figure 8 shows that the pattern is mostly decided by the tasks, where both heads shows similar pattern in the same task. In R.Number, we are able to observe a strong signal between token 10000 and 15000, which is the position of the inserted number string, indicating that Locret can identify the potentially answer-related parts by giving high predicted values of CIS. In M.Find, we can observe the StreamingLLM (Xiao et al., 2024c) pattern, where the tokens at the beginning of the sequence are always important. This is also mentioned as the $\Lambda$-pattern in MInference. We can also discover the vertical lines in the middle of the sequence. This pattern is also approached by MInference (Jiang et al., 2024a) by the pattern “vertical-and-slash”. In E.MC, H2O (Zhang et al., 2024d) and ScissorHands (Liu et al., 2024a) pattern can be observed, following the assumption that if a token is activated at some point, it will continue to be activated in the consequencing process. Noticing that the vertical lines always come in groups, which is the fundament of InfLLM (Xiao et al., 2024a) retrieving blocks to calculate. The comparison between two heads also shows that different heads exhibits different features. Head 22 of layer 11 shows stronger vertical lines at some point, where retained pattern of head 14 layer 11 is more even. Head 14 of layer 11 also gives stronger signal to the initial tokens, where this effect is less strong in head 22 layer 11. We also conduct experiments to investigate the patterns across layers. In Figure 8, we show that the pattern variance of the same head in different layers can be large. In shallow layers, e.g. layer 1 and 5, the retained cache units appears to be periodical and semantic independent. However, in middle layers, e.g. layer 13 and 17, the position of the inserted number string is strongly highlighted, indicating that semantic takes over to be the dominant factor. In the deepest layers, e.g. 21, 25 and 29, the highlighted vertical line at the position of the inserted string becomes more accurate.
@@
Appendix P The Loss Dynamics of Locret